Pat is paid time and a half for each hour he works over 36 hours in a week. last week he worked 42 hours for a total of $ 472.50. what is his normal hourly rate.
thank tou

Answer by stanbon(75887)

Pat is paid time and a half for each hour he works over 36 hours in a week. last week he worked 42 hours for a total of $ 472.50. what is his normal hourly rate.
-------
36 regular hrs; 6 overtime hrs
----
Let his normal rate be "x":
Then his overtime rate is (3/2)x
-------------------------------
36x + 6((3/2)x) = 472.50
------
36x + 9x = 472.50
45x = 472.50
x = $10.50 (normal hourly rate)
(3/2)x = (3/2)(10.50) = $15.75
==================================
Cheers,
Stan H.
=================